OpenStudy (vannah):

Im in 7th grade but i do college courses and need some help with this question,simplify 8 56/6

OpenStudy (anonymous):

is that 8 x (56/6)

OpenStudy (chaise):

8 56/6 Multiply the denominator by the number outside the fraction, plus the numerator then finally divide by the denominator. 104/6 = 52/3

OpenStudy (vannah):

the whole number is 8 56 is the numerator and 6 is the denominator

OpenStudy (anonymous):

428/3

OpenStudy (anonymous):

common factor of 2 for numerator and denominator

OpenStudy (vannah):

@snaef99 could you plsz show me how you got that

OpenStudy (anonymous):

divide the numerator and denominator by 2

OpenStudy (anonymous):

i think its 856/6

OpenStudy (anonymous):

not sure why that space is there

OpenStudy (vannah):

8 is the whole numer snaef99

OpenStudy (chaise):

The answer is what I first said it was. -.-

OpenStudy (anonymous):

oh!, you said "the whole number is 8 56 is the numerator and 6 is the denominator " that through me off

OpenStudy (vannah):

yheaa

OpenStudy (vannah):

\[8 \left(\begin{matrix}56 \\ 6\end{matrix}\right)\]

OpenStudy (chaise):

\[8\frac{56}{6}\]